# Faster Neighborhood Attention: Reducing the $\mathcal{O}(n^{2})$ Cost of Self Attention at the Threadblock Level

### v2

@@ 4 Experiments
NAT-M 20 2.7 2975 2660 ( $\boldsymbol{\downarrow}$11 % ) 3575 ( $\boldsymbol{\uparrow}$ 20 % ) 81.8
DiNAT-M 20 2.7 2672 2548 ( $\boldsymbol{\downarrow}$5 % ) 3781 ( $\boldsymbol{\uparrow}$ 42 % ) 81.8
NAT-T 28 4.3 2167 1939 ( $\boldsymbol{\downarrow}$11 % ) 2661 ( $\boldsymbol{\uparrow}$ 23 % ) 83.2
DiNAT-T 28 4.3 1910 1845 ( $\boldsymbol{\downarrow}$3 % ) 2818 ( $\boldsymbol{\uparrow}$ 48 % ) 82.7
NAT-S 51 7.8 1457 1309 ( $\boldsymbol{\downarrow}$10 % ) 1804 ( $\boldsymbol{\uparrow}$ 24 % ) 83.7
DiNAT-S 51 7.8 1360 1313 ( $\boldsymbol{\downarrow}$3 % ) 2070 ( $\boldsymbol{\uparrow}$ 52 % ) 83.8
NAT-B 90 13.7 1110 997 ( $\boldsymbol{\downarrow}$10 % ) 1384 ( $\boldsymbol{\uparrow}$ 25 % ) 84.3
DiNAT-B 90 13.7 982 950 ( $\boldsymbol{\downarrow}$3 % ) 1463 ( $\boldsymbol{\uparrow}$ 49 % ) 84.4
DiNAT-L 200 30.6 669 647 ( $\boldsymbol{\downarrow}$3 % ) 1006 ( $\boldsymbol{\uparrow}$ 50 % ) 86.6
DiNAT-L(384 × 384) 200 92.4 153 134 ( $\boldsymbol{\downarrow}$12 % ) 301 ( $\boldsymbol{\uparrow}$ 97 % ) 87.5
Table 5: Model-level throughput changes when using our proposed GEMM-based and fused kernels in ImageNet classification. Hierarchical vision transformers NAT and DiNAT can see between 25% to 106% improvement in FP16 throughput on an A100 (batch size 128) with our proposed fused kernel. Suffering from the memory alignment issue, our half precision GEMM kernels usually result in a much smaller improvement over naive kernels, particularly the tiled variants. The same measurements with FP32 precision are presented in Tab. 6.
@@
NAT-M 20 2.7 2416 2481 ( $\boldsymbol{\uparrow}$ 3 % ) 2408 ( 0 % ) 81.8
DiNAT-M 20 2.7 2217 2364 ( $\boldsymbol{\uparrow}$ 7 % ) 2690 ( $\boldsymbol{\uparrow}$ 21 % ) 81.8
NAT-T 28 4.3 1739 1802 ( $\boldsymbol{\uparrow}$ 4 % ) 1776 ( $\boldsymbol{\uparrow}$ 2 % ) 83.2
DiNAT-T 28 4.3 1591 1706 ( $\boldsymbol{\uparrow}$ 7 % ) 1988 ( $\boldsymbol{\uparrow}$ 25 % ) 82.7
NAT-S 51 7.8 1160 1199 ( $\boldsymbol{\uparrow}$ 3 % ) 1185 ( $\boldsymbol{\uparrow}$ 2 % ) 83.7
DiNAT-S 51 7.8 1102 1183 ( $\boldsymbol{\uparrow}$ 7 % ) 1389 ( $\boldsymbol{\uparrow}$ 26 % ) 83.8
NAT-B 90 13.7 867 897 ( $\boldsymbol{\uparrow}$ 3 % ) 890 ( $\boldsymbol{\uparrow}$ 3 % ) 84.3
DiNAT-B 90 13.7 795 851 ( $\boldsymbol{\uparrow}$ 7 % ) 995 ( $\boldsymbol{\uparrow}$ 25 % ) 84.4
DiNAT-L 200 30.6 506 540 ( $\boldsymbol{\uparrow}$ 7 % ) 629 ( $\boldsymbol{\uparrow}$ 24 % ) 86.6
DiNAT-L(384 × 384) 200 92.4 116 115 ( $\boldsymbol{\downarrow}$1 % ) 162 ( $\boldsymbol{\uparrow}$ 40 % ) 87.5
@@
FFHQ (256 × 256) 48.9 M 36.7 40.6 ( $\boldsymbol{\uparrow}$ 11 % ) 44.2 ( $\boldsymbol{\uparrow}$ 21 % ) 2.05
FFHQ (1024 × 1024) 49.4 M 8.2 8.5 ( $\boldsymbol{\uparrow}$ 3 % ) 10.4 ( $\boldsymbol{\uparrow}$ 26 % ) 4.17
5 Conclusion
In this work, we formulated the neighborhood attention problem, and by extension the sliding window attention problem, in both single-rank (1-D) and higher-rank spaces (2-D and 3-D), which are inherently GEMV problems as GEMM problems. Through this finding, we implemented extensible GEMM-based and fused CUDA kernels that implement neighborhood attention, which can significantly improve upon existing kernels in the $\mathcal{N}ATTEN$ project. These kernels will not only speed up previously-proposed models based on neighborhood attention, but can also significantly enhance ongoing research efforts in this direction. In addition, our fused kernels are at the same time the most flexible in terms of parameterization (support for varying window size, dilation factor, and causal masking across different ranks) and reduce the memory footprint of unfused (BMM-style) implementations. We however note that this is an ongoing effort, which does not yet include support for back-propagation, but at the same time we see no obstacles in extending this implementation to support backward passes, as well as additional features.
Acknowledgements. We would like to thank NVIDIA and members of the CUTLASS project [18], in particular Haicheng Wu, for his valuable feedback and comments which led to the creation of GEMM-based NA. We also thank Meta xFormers team for developing FMHA, which is what our fused neighborhood attention kernels are based on. A. H. thanks Bing Xu, Hao Lu, and Michael Iovine for the invaluable learning experience while interning at HippoML. A. H. also thanks Kai Wang for his feedback on the paper. This research was supported in part by the Intelligence Advanced Research Projects Activity (IARPA) under Contract No. 2022-21102100004, and cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA.

### v3

@@ Appendix B Additional experiments
NAT-M 20 2.7 2975 2660 ( $\boldsymbol{\downarrow}$11 % ) 3742 ( $\boldsymbol{\uparrow}$ 26 % ) 81.8
DiNAT-M 20 2.7 2672 2548 ( $\boldsymbol{\downarrow}$5 % ) 3930 ( $\boldsymbol{\uparrow}$ 47 % ) 81.8
DiNATs-T 28 4.5 2850 2504 ( $\boldsymbol{\downarrow}$12 % ) 3847 ( $\boldsymbol{\uparrow}$ 35 % ) 81.8
NAT-T 28 4.3 2167 1939 ( $\boldsymbol{\downarrow}$11 % ) 2772 ( $\boldsymbol{\uparrow}$ 28 % ) 83.2
DiNAT-T 28 4.3 1910 1845 ( $\boldsymbol{\downarrow}$3 % ) 2909 ( $\boldsymbol{\uparrow}$ 52 % ) 82.7
DiNATs-S 50 8.7 1800 1571 ( $\boldsymbol{\downarrow}$13 % ) 2445 ( $\boldsymbol{\uparrow}$ 36 % ) 83.5
NAT-S 51 7.8 1457 1309 ( $\boldsymbol{\downarrow}$10 % ) 1879 ( $\boldsymbol{\uparrow}$ 29 % ) 83.7
DiNAT-S 51 7.8 1360 1313 ( $\boldsymbol{\downarrow}$3 % ) 2145 ( $\boldsymbol{\uparrow}$ 58 % ) 83.8
DiNATs-B 88 15.4 1351 1178 ( $\boldsymbol{\downarrow}$13 % ) 1837 ( $\boldsymbol{\uparrow}$ 36 % ) 83.8
NAT-B 90 13.7 1110 997 ( $\boldsymbol{\downarrow}$10 % ) 1448 ( $\boldsymbol{\uparrow}$ 30 % ) 84.3
DiNAT-B 90 13.7 982 950 ( $\boldsymbol{\downarrow}$3 % ) 1517 ( $\boldsymbol{\uparrow}$ 54 % ) 84.4
DiNATs-L 197 34.5 846 744 ( $\boldsymbol{\downarrow}$12 % ) 1119 ( $\boldsymbol{\uparrow}$ 32 % ) 86.5
DiNAT-L 200 30.6 669 647 ( $\boldsymbol{\downarrow}$3 % ) 1042 ( $\boldsymbol{\uparrow}$ 56 % ) 86.6
DiNATs-L(384 × 384) 197 101.5 295 239 ( $\boldsymbol{\downarrow}$19 % ) 391 ( $\boldsymbol{\uparrow}$ 33 % ) 87.4
DiNAT-L(384 × 384) 200 92.4 153 134 ( $\boldsymbol{\downarrow}$12 % ) 312 ( $\boldsymbol{\uparrow}$ 104 % ) 87.5
@@
NAT-M 20 2.7 2416 2481 ( $\boldsymbol{\uparrow}$ 3 % ) 2658 ( $\boldsymbol{\uparrow}$ 10 % ) 81.8
DiNAT-M 20 2.7 2217 2364 ( $\boldsymbol{\uparrow}$ 7 % ) 2905 ( $\boldsymbol{\uparrow}$ 31 % ) 81.8
DiNATs-T 28 4.5 2270 2255 ( $\boldsymbol{\downarrow}$1 % ) 2771 ( $\boldsymbol{\uparrow}$ 22 % ) 81.8
NAT-T 28 4.3 1739 1802 ( $\boldsymbol{\uparrow}$ 4 % ) 1942 ( $\boldsymbol{\uparrow}$ 12 % ) 83.2
DiNAT-T 28 4.3 1591 1706 ( $\boldsymbol{\uparrow}$ 7 % ) 2123 ( $\boldsymbol{\uparrow}$ 33 % ) 82.7
DiNATs-S 50 8.7 1403 1393 ( $\boldsymbol{\downarrow}$1 % ) 1717 ( $\boldsymbol{\uparrow}$ 22 % ) 83.5
NAT-S 51 7.8 1160 1199 ( $\boldsymbol{\uparrow}$ 3 % ) 1293 ( $\boldsymbol{\uparrow}$ 11 % ) 83.7
DiNAT-S 51 7.8 1102 1183 ( $\boldsymbol{\uparrow}$ 7 % ) 1490 ( $\boldsymbol{\uparrow}$ 35 % ) 83.8
DiNATs-B 88 15.4 1020 1009 ( $\boldsymbol{\downarrow}$1 % ) 1240 ( $\boldsymbol{\uparrow}$ 22 % ) 83.8
NAT-B 90 13.7 867 897 ( $\boldsymbol{\uparrow}$ 3 % ) 966 ( $\boldsymbol{\uparrow}$ 11 % ) 84.3
DiNAT-B 90 13.7 795 851 ( $\boldsymbol{\uparrow}$ 7 % ) 1059 ( $\boldsymbol{\uparrow}$ 33 % ) 84.4
DiNATs-L 197 34.5 609 601 ( $\boldsymbol{\downarrow}$1 % ) 721 ( $\boldsymbol{\uparrow}$ 18 % ) 86.5
DiNAT-L 200 30.6 506 540 ( $\boldsymbol{\uparrow}$ 7 % ) 669 ( $\boldsymbol{\uparrow}$ 32 % ) 86.6
DiNATs-L(384 × 384) 197 101.5 211 193 ( $\boldsymbol{\downarrow}$9 % ) 245 ( $\boldsymbol{\uparrow}$ 16 % ) 87.4
DiNAT-L(384 × 384) 200 92.4 116 115 ( $\boldsymbol{\downarrow}$1 % ) 179 ( $\boldsymbol{\uparrow}$ 54 % ) 87.5
@@
FFHQ (256 × 256) 48.9 M 36.7 40.6 ( $\boldsymbol{\uparrow}$ 11 % ) 45.5 ( $\boldsymbol{\uparrow}$ 24 % ) 2.05
FFHQ (1024 × 1024) 49.4 M 8.2 8.5 ( $\boldsymbol{\uparrow}$ 3 % ) 11.5 ( $\boldsymbol{\uparrow}$ 40 % ) 4.17
Finally, we also attempted to estimate improvements in training time compared to our baseline. As suggested by our earlier findings regarding the limit of our GEMM-based implementation in the backward pass, we do not see any improvement in training time compared to the naive baseline. However, we find that our fused kernels deliver on the promise of improved half precision training time. We present our estimates in Tab. 8, which are based on measurements from training NAT [9] and DiNAT [8] variants according to their original specifications. We ran each model for 1 warmup epoch, and 1 benchmark epoch, the average throughput of which is used to estimate training time for 300 epochs.
Table 8: Training time improvement when using fused neighborhood attention kernels. We ran each of the classification models based on neighborhood attention for one warmup epoch and one benchmark epoch, all with half precision (the typical training scenario), and report the estimated training time. Note that these numbers exclude positional biases, as our fused backward kernel does not support it.
Model # of FLOPs Training time estimate
Params Naive GEMM Fused
(M) (G) (hours)
NAT-M 20 2.7 19.4 20.4 ( $\boldsymbol{\downarrow}$5 % ) 16.6 ( $\boldsymbol{\uparrow}$ 17 % )
DiNAT-M 20 2.7 20.4 21.2 ( $\boldsymbol{\downarrow}$4 % ) 17.4 ( $\boldsymbol{\uparrow}$ 17 % )
DiNATs-T 28 4.5 21.1 22.0 ( $\boldsymbol{\downarrow}$4 % ) 17.4 ( $\boldsymbol{\uparrow}$ 21 % )
NAT-T 28 4.3 26.5 28.2 ( $\boldsymbol{\downarrow}$6 % ) 24.0 ( $\boldsymbol{\uparrow}$ 10 % )
DiNAT-T 28 4.3 27.4 28.5 ( $\boldsymbol{\downarrow}$4 % ) 21.9 ( $\boldsymbol{\uparrow}$ 25 % )
DiNATs-S 50 8.7 33.3 33.2 ( 0 % ) 25.1 ( $\boldsymbol{\uparrow}$ 33 % )
NAT-S 51 7.8 39.2 41.8 ( $\boldsymbol{\downarrow}$6 % ) 33.7 ( $\boldsymbol{\uparrow}$ 16 % )
DiNAT-S 51 7.8 38.0 40.1 ( $\boldsymbol{\downarrow}$5 % ) 30.8 ( $\boldsymbol{\uparrow}$ 23 % )
DiNATs-B 88 15.4 45.4 46.1 ( $\boldsymbol{\downarrow}$2 % ) 32.6 ( $\boldsymbol{\uparrow}$ 39 % )
NAT-B 90 13.7 51.1 54.6 ( $\boldsymbol{\downarrow}$6 % ) 47.7 ( $\boldsymbol{\uparrow}$ 7 % )
DiNAT-B 90 13.7 54.4 56.0 ( $\boldsymbol{\downarrow}$3 % ) 41.0 ( $\boldsymbol{\uparrow}$ 33 % )